At Mc. Donald's, the barbecue sauce packet has five fewer calories than the sweet n sour sauce packet. Ten barbecue sauce packets have the same amount of calories as nine sweet n sour sauce packets. how many calories does each have?
Answer by robertb(5830)

b = s - 5,
10b = 9s.
Solve the system. By substitution,
10(s-5) = 9s,
10s - 50 = 9s,
s = 50, the # of calories in 1 packet of the barbecue sauce;
s - 5 = 45, the # of calories in 1 packet of the sweet n sour sauce packet.